# Approximate independence of permutation mixtures

### v2

@@ Appendix C Deferred proofs
Here (a) follows from convexity and $P_{X_{\backslash i}|X_{i}}=\mathbb{E}_{\vartheta_{i}^{\prime}|X_{i}}\left[P_{X_{\backslash i}|\vartheta_{i}^{\prime}}\right]$, where $\vartheta_{i}^{\prime}$ is an independent copy of $\vartheta_{i}$ given $X_{i}$. For (b), note that $P_{X_{\backslash i}|\vartheta_{i}}$ is the permutation mixture based on $\{P_{\theta_{1}},\dots,P_{\theta_{n}}\}\backslash\{P_{\vartheta_{i}}\}$, so that Theorem 1.6 gives the claimed upper bound. Plugging this upper bound into (45) and using $n\geq 2$ leads to the claimed corollary.
C.5 Proof of Lemma 4.1
@@
=(26)${=}\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{n}(nA_{\pi(i),\pi^{\prime}(i)})\right]=\frac{n^{n}}{n!}\mathrm{Perm}(A),$
@@
As for the eigenstructure of $A$, the same leading eigenvalue/eigenvector holds for any doubly stochastic matrix. The only non-trivial property is the spectral gap. To this end, note that the Laplacian $L=I-A$ satisfies that
@@
$x^{\top}Lx\geq\frac{1}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}\sum_{1\leq i<j\leq n}(x_{i}-x_{j})^{2}=\frac{1}{2\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}\sum_{i,j=1}^{n}(x_{i}-x_{j})^{2}=\frac{1}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})},$
@@
$\mathbb{E}\left[\prod_{i=1}^{m}|(Pz)_{i}|^{2}\right]$ $=\mathbb{E}\left[\prod_{i=1}^{m}\left|\sum_{j=1}^{n}P_{ij}z_{j}\right|^{2}\right]$
$=\mathbb{E}\left[\prod_{i=1}^{m}\left(\sum_{j=1}^{n}P_{ij}z_{j}\right)\left(\sum_{j=1}^{n}P_{ij}\bar{z}_{j}\right)\right]$
=$\overset{\rm(a)}{=}\sum_{\pi\in S_{m}}\prod_{i=1}^{m}\mathbb{E}\left[\left(\sum_{j=1}^{n}P_{ij}z_{j}\right)\left(\sum_{j=1}^{n}P_{\pi(i)j}\bar{z}_{j}\right)\right]$
$=\sum_{\pi\in S_{m}}\prod_{i=1}^{m}\left(\sum_{j=1}^{n}P_{ij}P_{\pi(i)j}\right)$
@@
C.8 Proof of Lemma 4.6
Let $v=(v_{1},\cdots,v_{n}):=\widetilde{P}z\sim{\mathcal{C}}{\mathcal{N}}(0,\widetilde{P}\widetilde{P}^{\top})$, where
@@
where $\lambda_{1}=1$ and $u_{1}={\bf 1}/\sqrt{n}$ are the leading eigenvalue and eigenvector of $A$, respectively (cf. Lemma 4.2), and $\overline{A}$ is the centered version of $A$ defined above (33). Consequently,
$\mathbb{E}\left[|e_{\ell}(v)|^{2}\right]$ $=\mathbb{E}\left[\left|\sum_{S\subseteq[n]:|S|=\ell}\prod_{i\in S}v_{i}\right|^{2}\right]=\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=\ell\end{subarray}}\mathbb{E}\left[\prod_{i\in S}v_{i}\prod_{j\in S^{\prime}}\bar{v}_{j}\right]$
=$\overset{\rm(a)}{=}\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=\ell\end{subarray}}\sum_{\pi:S\leftrightarrow S^{\prime}}\prod_{i\in S}\mathbb{E}\left[v_{i}\bar{v}_{\pi(i)}\right]$
@@
where (a) is due to Isserlis’ theorem [Iss18] and the observation $\mathbb{E}[v_{i}v_{j}]=\mathbb{E}[\bar{v}_{i}\bar{v}_{j}]=0$ for all $i,j\in[n]$ (here $\pi:S\leftrightarrow S^{\prime}$ denotes that $\pi$ is a bijection between $S$ and $S^{\prime}$), and (b) follows from $v\sim{\mathcal{C}}{\mathcal{N}}(0,\overline{A})$. Then the claimed result is a direct consequence of the identity between $S_{\ell}$ and $T_{\ell}$ in Lemma A.1.
C.9 Proof of Lemma 4.7
${\sf TV}(\mathbb{P}_{n},\mathbb{Q}_{n})^{2}=\left(\frac{1}{2}\int|\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{Q}_{n}|\right)^{2}\leq\frac{1}{4}\int\frac{(\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{Q}_{n})^{2}}{\mathrm{d}\mathbb{R}_{n}}.$
Choosing $\mathbb{R}_{n}=\overline{P}^{\otimes n}$ completes the proof of the first claim. The likelihood ratio between the signed measure $\mathbb{P}_{n}-\mathbb{Q}_{n}$ and the product measure $\overline{P}^{\otimes n}$ is computed as
$\frac{\mathrm{d}(\mathbb{P}_{n}-\mathbb{Q}_{n})}{\mathrm{d}\overline{P}^{\otimes n}}(X)$ =$\overset{\rm(a)}{=}\mathbb{E}_{I\sim\text{\rm Unif}([n]),\pi:[n]\backslash\{I\}\leftrightarrow\{2,\ldots,n\}}\left[\frac{\mathrm{d}(P_{1}-Q_{1})}{\mathrm{d}\overline{P}}(X_{I})\cdot\prod_{i\neq I}\frac{\mathrm{d}P_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\right]$
=$\overset{\rm(b)}{=}\mathbb{E}_{I\sim\text{\rm Unif}([n]),\pi:[n]\backslash\{I\}\leftrightarrow\{2,\ldots,n\}}\left[\frac{\mathrm{d}(P_{1}-Q_{1})}{\mathrm{d}\overline{P}}(X_{I})\cdot\sum_{T\subseteq[n]\backslash\{I\}}\prod_{i\in T}\frac{\mathrm{d}(P_{\pi(i)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{i})\right]$
=$\overset{\rm(c)}{=}\sum_{S\subseteq[n]:|S|\geq 1}\left(\frac{1}{n}\sum_{i\in S}\frac{\mathrm{d}(P_{1}-Q_{1})}{\mathrm{d}\overline{P}}(X_{i})\cdot\mathbb{E}_{\pi:[n]\backslash\{i\}\leftrightarrow\{2,\ldots,n\}}\prod_{j\in S\backslash\{i\}}\frac{\mathrm{d}(P_{\pi(j)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{j})\right)$
$=:\sum_{S\subseteq[n]:|S|\geq 1}f_{S}(X),$
where (a) decomposes the random permutation over $S_{n}$ into a random choice of $I=\pi^{-1}(1)$ and a random bijection between the rest (i.e. $[n]\backslash\{i\}$ and $\{2,\cdots,n\}$), and (b) is the identity $\prod_{i=1}^{n}(1+x_{i})=\sum_{T\subseteq[n]}\prod_{i\in T}x_{i}$. In (c), we expand the expectation over $I\sim\text{\rm Unif}([n])$ and swap the sum as $\sum_{i=1}^{n}\sum_{T\subseteq[n]\backslash\{i\}}g_{i}(T)=\sum_{S\subseteq[n]:|S|\geq 1}\sum_{i\in S}g_{i}(S\backslash\{i\})$.
$g_{S,i,\pi}(X)=\frac{\mathrm{d}(P_{1}-Q_{1})}{\mathrm{d}\overline{P}}(X_{i})\prod_{j\in S\backslash\{i\}}\frac{\mathrm{d}(P_{\pi(j)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{j})$
$\int\frac{(\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{Q}_{n})^{2}}{\mathrm{d}\overline{P}^{\otimes n}}=\mathbb{E}_{\overline{P}^{\otimes n}}\left[\left(\sum_{S\subseteq[n]:|S|\geq 1}f_{S}(X)\right)^{2}\right]=\sum_{S\subseteq[n]:|S|\geq 1}\mathbb{E}_{\overline{P}^{\otimes n}}\left[f_{S}(X)^{2}\right]$
=$\overset{\rm(d)}{=}\sum_{\ell=1}^{n}\binom{n}{\ell}\mathbb{E}_{\overline{P}^{\otimes n}}\left[f_{[\ell]}(X)^{2}\right]$
$\overset{\rm(e)}{\leq}\sum_{\ell=1}^{n}\binom{n}{\ell}\frac{\ell}{n^{2}}\sum_{i=1}^{\ell}\mathbb{E}_{X_{1},\ldots,X_{\ell}\sim\overline{P}}\left[\left(\frac{\mathrm{d}(P_{1}-Q_{1})}{\mathrm{d}\overline{P}}(X_{i})\cdot\mathbb{E}_{\pi:[n]\backslash\{i\}\leftrightarrow\{2,\ldots,n\}}\prod_{j\in[\ell]\backslash\{i\}}\frac{\mathrm{d}(P_{\pi(j)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{j})\right)^{2}\right]$
=$\overset{\rm(f)}{=}\sum_{\ell=1}^{n}\binom{n}{\ell}\frac{\ell^{2}}{n^{2}}\int\frac{(\mathrm{d}P_{1}-\mathrm{d}Q_{1})^{2}}{\mathrm{d}\overline{P}}\cdot\mathbb{E}_{X_{2},\ldots,X_{\ell}\sim\overline{P}}\left[\left(\mathbb{E}_{\pi:\{2,\ldots,n\}\leftrightarrow\{2,\ldots,n\}}\prod_{j=2}^{\ell}\frac{\mathrm{d}(P_{\pi(j)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{j})\right)^{2}\right],$
$\int\frac{(\mathrm{d}P_{1}-\mathrm{d}Q_{1})^{2}}{\mathrm{d}\overline{P}}$ $\leq 2\int\frac{(\mathrm{d}P_{1}-\mathrm{d}\overline{P})^{2}}{\mathrm{d}\overline{P}}+2\int\frac{(\mathrm{d}Q_{1}-\mathrm{d}\overline{P})^{2}}{\mathrm{d}\overline{P}}$
$=2(\chi^{2}(P_{1}\|\overline{P})+\chi^{2}(Q_{1}\|\overline{P}))\leq 4{\mathsf{D}}_{\chi^{2}}({\mathcal{P}}),$
and
$\mathbb{E}_{X_{2},\ldots,X_{\ell}\sim\overline{P}}\left[\left(\mathbb{E}_{\pi:\{2,\ldots,n\}\leftrightarrow\{2,\ldots,n\}}\prod_{j=2}^{\ell}\frac{\mathrm{d}(P_{\pi(j)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{j})\right)^{2}\right]=R_{\ell-1}=\frac{S_{\ell-1}}{\binom{n-1}{\ell-1}}$
by the definition of $R_{\ell}$ in (31) and the identity between $R_{\ell}$ and $S_{\ell}$ in Lemma A.1, where the matrix $A\in\mathbb{R}^{(n-1)\times(n-1)}$ is now constructed from $(P_{2},\ldots,P_{n})$. A combination of all the above then gives
${\sf TV}(\mathbb{P}_{n},\mathbb{Q}_{n})^{2}\leq\frac{1}{4}\sum_{\ell=1}^{n}\binom{n}{\ell}\frac{\ell^{2}}{n^{2}}\cdot 4{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})\frac{S_{\ell-1}}{\binom{n-1}{\ell-1}}=\frac{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n}\sum_{\ell=1}^{n}\ell S_{\ell-1},$
as claimed.
C.10 Proof of Lemma 5.1
@@
=$\overset{\rm(c)}{=}\left[1-\left(1-\frac{\frac{1}{n}\sum_{i=1}^{n}\mathrm{Var}_{P_{i}}(f)}{\mathrm{Var}_{\overline{P}}(f)}\right)^{2}\right]^{-\frac{1}{2}}-1,$
where (a) is due to the data-processing inequality, (b) is the lower-semicontinuity of the $\chi^{2}$ divergence (using arguments similar to [PW24, Theorem 4.9]), and (c) uses
@@
with $u=(u_{1},\dots,u_{n})$ being the unit-length eigenvector of $A$ associated with $\lambda_{2}(A)$. Clearly $\mathbb{E}_{\overline{P}}[f]=\sum_{j=1}^{n}u_{j}=0$ by the orthogonality of $u$ and $\bf 1$ in Lemma 4.2, and
@@
This shows that $S\geq\lambda_{2}(A)$ (in fact $S=\lambda_{2}(A)$ holds, but we do not need this upper bound).
@@
Here the last step follows from [PW24, Eqn. (7.35)]. Consequently, the second statement follows from
@@
where $I$ and $J$ are the identity and all-one matrices, respectively. By simple algebra, $A$ is doubly stochastic, and has eigenvalues $\lambda_{1}(A)=1,\lambda_{2}(A)=\cdots=\lambda_{m}(A)=1-\Delta$, and $\lambda_{m+1}(A)=\cdots=\lambda_{mn}(A)=0$. In particular, $\mathrm{Tr}(A)\leq m<1+{\mathsf{C}}$, and the spectral gap of $A$ is $\Delta$.

### v3

@@ Appendix C Deferred proofs
[…]
C.8 Proof of Lemma˜2.4
For the first inequality (11), the triangle inequality applied to the separable loss yields
[…]
$\leq\frac{1}{n}\sum_{i=1}^{n}\sqrt{\chi^{2}(\mathbb{P}_{\vartheta_{i},X}\|\mathbb{Q}_{\vartheta_{i},X})\mathrm{Var}_{\mathbb{Q}}[\ell(\vartheta_{i},\widehat{\vartheta}_{i})]}.$
[…]
Conditioned on $\vartheta_{i}=\theta_{j}$, it is clear that $\mathbb{P}_{X_{\backslash i}|\vartheta_{i}}$ is the (n1)(n-1)-dimensional permutation mixture generated by $\{P_{\theta_{k}}:k\neq j\}$, and $\mathbb{Q}_{X_{\backslash i}|\vartheta_{i}}$ is the i.i.d. distribution with marginal $\frac{1}{n}\sum_{i=1}^{n}P_{\theta_{i}}$. Defining an auxiliary distribution $\mathbb{Q}_{n-1}=(\frac{1}{n-1}\sum_{k\neq j}P_{\theta_{k}})^{\otimes(n-1)}$, then
$\frac{\mathrm{d}\mathbb{Q}_{X_{\backslash i}|\vartheta_{i}=\theta_{j}}}{\mathrm{d}\mathbb{Q}_{n-1}}\geq\left(\frac{n-1}{n}\right)^{n-1}\geq\frac{1}{e}$
almost surely. Consequently,
$\chi^{2}\left(\mathbb{P}_{X_{\backslash i}|\vartheta_{i}=\theta_{j}}\|\mathbb{Q}_{X_{\backslash i}|\vartheta_{i}=\theta_{j}}\right)+1$ $=\int\frac{(\mathrm{d}\mathbb{P}_{X_{\backslash i}|\vartheta_{i}=\theta_{j}})^{2}}{\mathrm{d}\mathbb{Q}_{X_{\backslash i}|\vartheta_{i}=\theta_{j}}}\leq e\int\frac{(\mathrm{d}\mathbb{P}_{X_{\backslash i}|\vartheta_{i}=\theta_{j}})^{2}}{\mathrm{d}\mathbb{Q}_{n-1}}$
$=e\left(\chi^{2}(\mathbb{P}_{X_{\backslash i}|\vartheta_{i}=\theta_{j}}\|\mathbb{Q}_{n-1})+1\right)\overset{\rm(a)}{\leq}e(\chi^{2}({\mathcal{P}})+1),$
where (a) notes that the i.i.d. distribution $\mathbb{Q}_{n-1}$ has the same one-dimensional marginal as $\mathbb{P}_{X_{\backslash i}|\vartheta_{i}=\theta_{j}}$. This establishes the inequality (11). The second inequality (12) follows from the concavity of $x\mapsto\sqrt{x}$, and the variance upper bound
[…]
C.9 Proof of Lemma˜2.5
We begin with a technical result. Let $P_{0},\dots,P_{n}\in{\mathcal{P}}$ be $n+1$ probability distributions, and for $i\in\{0,\dots,n\}$, define the permutation mixture $\mathbb{P}_{-i}$ as
[…]
Lemma C.2.
For $\overline{\mathbb{P}}=\frac{1}{n+1}\sum_{i=0}^{n}\mathbb{P}_{-i}$ and every $i\in\{0,\dots,n\}$ it holds that
In particular, for every $i,j\in\{0,\dots,n\}$ we have
The proof of Lemma˜C.2 is deferred to the end of this section. Under the quadratic loss, it is shown in [GR09] that the optimal simple and permutation invariant decision rules are
respectively, where the joint distribution of $(\vartheta,X)$ is given in the “postulated Bayes model” in the proof of Lemma˜2.3. In addition, the risk difference admits a simple form
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=\mathbb{E}\left[\left(\mathbb{E}\left[\vartheta_{1}|X_{1}\right]-\mathbb{E}\left[\vartheta_{1}|X\right]\right)^{2}\right].$
Since
$\mathbb{E}\left[\vartheta_{1}|X_{1}\right]-\mathbb{E}\left[\vartheta_{1}|X\right]=\sum_{i=1}^{n}\theta_{i}\left(P_{\vartheta_{1}=\theta_{i}|X_{1}}-P_{\vartheta_{1}=\theta_{i}|X}\right)=\sum_{i=1}^{n}\theta_{i}P_{\vartheta_{1}=\theta_{i}|X_{1}}\Bigg{(}1-\frac{P_{X_{2}^{n}|\vartheta_{1}=\theta_{i}}}{P_{X_{2}^{n}|X_{1}}}\Bigg{)}$
according to the Bayes rule, we have
$\overline{P}:=P_{X_{2}^{n}|X_{1}}=\sum_{i=1}^{n}w_{i}P_{i}.$
and therefore
To upper bound the largest eigenvalue of $\overline{A}$, note that Lemma˜C.2 implies that $H^{2}(P_{i},P_{j})\leq\frac{4{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n}$ for all $i,j\in[n]$. Consequently, following the spectral gap argument in Lemma˜5.2, it holds that
In addition,
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\Bigg{(}1+\chi^{2}\bigg{(}P_{\theta_{i}}\bigg{\|}\frac{1}{n}\sum_{j=1}^{n}P_{\theta_{j}}\bigg{)}\Bigg{)}\leq\frac{1+{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}{n}.$
A combination of the above inequalities shows that
as claimed.
Proof of Lemma˜C.2.
$P_{\pi}$ $:=P_{\pi(1)}\otimes\cdots\otimes P_{\pi(n)}$
$\chi^{2}(\overline{\mathbb{P}}\|\mathbb{P}_{-0})$ $=\chi^{2}\Bigg{(}\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\Bigg{[}\frac{1}{n+1}\sum_{j=0}^{n}P_{\pi,j}\Bigg{]}\bigg{\|}\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[P_{\pi}\right]\Bigg{)}$
$\leq\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\Bigg{[}\chi^{2}\Bigg{(}\frac{1}{n+1}\sum_{j=0}^{n}P_{\pi,j}\bigg{\|}P_{\pi}\Bigg{)}\Bigg{]}$
$\overset{\rm(a)}{\leq}\frac{n}{n+1}\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\Bigg{[}\chi^{2}\Bigg{(}\frac{1}{n}\sum_{j=1}^{n}P_{\pi,j}\bigg{\|}P_{\pi}\Bigg{)}\Bigg{]}.$
Consequently, for $j\neq j^{\prime}$,
and
Therefore, the second moment method gives
$\chi^{2}\Bigg{(}\frac{1}{n}\sum_{j=1}^{n}P_{\pi,j}\bigg{\|}P_{\pi}\Bigg{)}=\frac{1}{n^{2}}\sum_{j,j^{\prime}=1}^{n}\int\frac{\mathrm{d}P_{\pi,j}\mathrm{d}P_{\pi,j^{\prime}}}{\mathrm{d}P_{\pi}}-1\leq\frac{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n},$
and we obtain the first claim. The second claim follows from
where (b) is the triangle inequality of the Hellinger distance, and (c) uses $H^{2}(P,Q)\leq\chi^{2}(P\|Q)$. ∎
C.10 Proof of Lemma˜5.1
@@
[…]
@@
As for the eigenstructure of $A$, the same leading eigenvalue/eigenvector holds for any doubly stochastic matrix. The only non-trivial property is the spectral gap. To this end, note that the Laplacian L=IAL=I-A satisfies that
@@
$x^{\top}Lx\geq\frac{1}{n\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}\sum_{1\leq i<j\leq n}(x_{i}-x_{j})^{2}=\frac{1}{2n\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}\sum_{i,j=1}^{n}(x_{i}-x_{j})^{2}=\frac{1}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})},$
@@
$\mathbb{E}\left[\prod_{i=1}^{m}|(Pz)_{i}|^{2}\right]$ $=\mathbb{E}\Bigg{[}\prod_{i=1}^{m}\Big{|}\sum_{j=1}^{n}P_{ij}z_{j}\Big{|}^{2}\Bigg{]}$
$=\mathbb{E}\Bigg{[}\prod_{i=1}^{m}\Bigg{(}\sum_{j=1}^{n}P_{ij}z_{j}\Bigg{)}\Bigg{(}\sum_{j=1}^{n}P_{ij}\bar{z}_{j}\Bigg{)}\Bigg{]}$
$\overset{\rm(a)}{=}\sum_{\pi\in S_{m}}\prod_{i=1}^{m}\mathbb{E}\Bigg{[}\Bigg{(}\sum_{j=1}^{n}P_{ij}z_{j}\Bigg{)}\Bigg{(}\sum_{j=1}^{n}P_{\pi(i)j}\bar{z}_{j}\Bigg{)}\Bigg{]}$
$=\sum_{\pi\in S_{m}}\prod_{i=1}^{m}\Bigg{(}\sum_{j=1}^{n}P_{ij}P_{\pi(i)j}\Bigg{)}$
@@
C.13 Proof of Lemma˜5.6
${\sf TV}(\mathbb{P}_{n},\mathbb{P}_{n}^{\prime})^{2}=\left(\frac{1}{2}\int|\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{P}_{n}^{\prime}|\right)^{2}\leq\frac{1}{4}\int\frac{(\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{P}_{n}^{\prime})^{2}}{\mathrm{d}\mathbb{R}_{n}}.$
Choosing $\mathbb{R}_{n}=\overline{P}^{\otimes n}$ completes the proof of the first claim. The likelihood ratio between the signed measure $\mathbb{P}_{n}-\mathbb{P}_{n}^{\prime}$ and the product measure $\overline{P}^{\otimes n}$ is computed as
$\frac{\mathrm{d}(\mathbb{P}_{n}-\mathbb{P}_{n}^{\prime})}{\mathrm{d}\overline{P}^{\otimes n}}(X)$ $\overset{\rm(a)}{=}\mathbb{E}_{I\sim\text{\rm Unif}([n]),\pi:[n]\backslash\{I\}\leftrightarrow\{2,\ldots,n\}}\Bigg{[}\frac{\mathrm{d}(P_{1}-P_{1}^{\prime})}{\mathrm{d}\overline{P}}(X_{I})\cdot\prod_{i\neq I}\frac{\mathrm{d}P_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\Bigg{]}$
$\overset{\rm(b)}{=}\mathbb{E}_{I\sim\text{\rm Unif}([n]),\pi:[n]\backslash\{I\}\leftrightarrow\{2,\ldots,n\}}\Bigg{[}\frac{\mathrm{d}(P_{1}-P_{1}^{\prime})}{\mathrm{d}\overline{P}}(X_{I})\cdot\sum_{T\subseteq[n]\backslash\{I\}}\prod_{i\in T}\frac{\mathrm{d}(P_{\pi(i)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{i})\Bigg{]}$
$\overset{\rm(c)}{=}\sum_{S\subseteq[n]:|S|\geq 1}\Bigg{(}\frac{1}{n}\sum_{i\in S}\frac{\mathrm{d}(P_{1}-P_{1}^{\prime})}{\mathrm{d}\overline{P}}(X_{i})\cdot\mathbb{E}_{\pi:[n]\backslash\{i\}\leftrightarrow\{2,\ldots,n\}}\prod_{j\in S\backslash\{i\}}\frac{\mathrm{d}(P_{\pi(j)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{j})\Bigg{)}$
$=:\sum_{S\subseteq[n]:|S|\geq 1}f_{S}(X),$
[…]
$g_{S,i,\pi}(X)=\frac{\mathrm{d}(P_{1}-P_{1}^{\prime})}{\mathrm{d}\overline{P}}(X_{i})\prod_{j\in S\backslash\{i\}}\frac{\mathrm{d}(P_{\pi(j)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{j})$
$\int\frac{(\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{P}_{n}^{\prime})^{2}}{\mathrm{d}\overline{P}^{\otimes n}}=\mathbb{E}_{\overline{P}^{\otimes n}}\Bigg{[}\Bigg{(}\sum_{S\subseteq[n]:|S|\geq 1}f_{S}(X)\Bigg{)}^{2}\Bigg{]}=\sum_{S\subseteq[n]:|S|\geq 1}\mathbb{E}_{\overline{P}^{\otimes n}}\left[f_{S}(X)^{2}\right]$
$\overset{\rm(d)}{=}\sum_{\ell=1}^{n}\binom{n}{\ell}\mathbb{E}_{\overline{P}^{\otimes n}}\left[f_{[\ell]}(X)^{2}\right]$
[…]
$\overset{\rm(f)}{=}\sum_{\ell=1}^{n}\binom{n}{\ell}\frac{\ell^{2}}{n^{2}}\int\frac{(\mathrm{d}P_{1}-\mathrm{d}P_{1}^{\prime})^{2}}{\mathrm{d}\overline{P}}\cdot\mathbb{E}_{X_{2},\ldots,X_{\ell}\sim\overline{P}}\Bigg{[}\Bigg{(}\mathbb{E}_{\pi:\{2,\ldots,n\}\leftrightarrow\{2,\ldots,n\}}\prod_{j=2}^{\ell}\frac{\mathrm{d}(P_{\pi(j)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{j})\Bigg{)}^{2}\Bigg{]},$
$\int\frac{(\mathrm{d}P_{1}-\mathrm{d}P_{1}^{\prime})^{2}}{\mathrm{d}\overline{P}}$ $\leq 2\int\frac{(\mathrm{d}P_{1}-\mathrm{d}\overline{P})^{2}}{\mathrm{d}\overline{P}}+2\int\frac{(\mathrm{d}P_{1}^{\prime}-\mathrm{d}\overline{P})^{2}}{\mathrm{d}\overline{P}}$
$=2(\chi^{2}(P_{1}\|\overline{P})+\chi^{2}(P_{1}^{\prime}\|\overline{P}))\leq 4{\mathsf{D}}_{\chi^{2}}({\mathcal{P}}),$
and
$\mathbb{E}_{X_{2},\ldots,X_{\ell}\sim\overline{P}}\Bigg{[}\Bigg{(}\mathbb{E}_{\pi:\{2,\ldots,n\}\leftrightarrow\{2,\ldots,n\}}\prod_{j=2}^{\ell}\frac{\mathrm{d}(P_{\pi(j)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{j})\Bigg{)}^{2}\Bigg{]}=R_{\ell-1}=\frac{S_{\ell-1}}{\binom{n-1}{\ell-1}}$
by the definition of $R_{\ell}$ in (34) and the identity between $R_{\ell}$ and $S_{\ell}$ in Lemma˜A.1, where the matrix $A\in\mathbb{R}^{(n-1)\times(n-1)}$ is now constructed from $(P_{2},\ldots,P_{n})$. A combination of all the above then gives
${\sf TV}(\mathbb{P}_{n},\mathbb{P}_{n}^{\prime})^{2}\leq\frac{1}{4}\sum_{\ell=1}^{n}\binom{n}{\ell}\frac{\ell^{2}}{n^{2}}\cdot 4{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})\frac{S_{\ell-1}}{\binom{n-1}{\ell-1}}=\frac{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n}\sum_{\ell=1}^{n}\ell S_{\ell-1},$
as claimed.
C.14 Proof of Lemma˜5.7
[…]
@@
where $\lambda_{1}=1$ and $u_{1}={\bf 1}/\sqrt{n}$ are the leading eigenvalue and eigenvector of $A$, respectively (cf. Lemma˜5.2), and $\overline{A}$ is the centered version of $A$ defined above (36). Consequently,
$\mathbb{E}\left[|e_{\ell}(v)|^{2}\right]$ $=\mathbb{E}\Bigg{[}\Big{|}\sum_{S\subseteq[n]:|S|=\ell}\prod_{i\in S}v_{i}\Big{|}^{2}\Bigg{]}=\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=\ell\end{subarray}}\mathbb{E}\Bigg{[}\prod_{i\in S}v_{i}\prod_{j\in S^{\prime}}\bar{v}_{j}\Bigg{]}$
$\overset{\rm(a)}{=}\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=\ell\end{subarray}}\sum_{\pi:S\leftrightarrow S^{\prime}}\prod_{i\in S}\mathbb{E}\left[v_{i}\bar{v}_{\pi(i)}\right]$
@@
where (a) is due to Isserlis’ theorem [Iss18] and the observation $\mathbb{E}[v_{i}v_{j}]=\mathbb{E}[\bar{v}_{i}\bar{v}_{j}]=0$ for all $i,j\in[n]$ (here $\pi:S\leftrightarrow S^{\prime}$ denotes that $\pi$ is a bijection between $S$ and $S^{\prime}$), and (b) follows from $v\sim{\mathcal{C}}{\mathcal{N}}(0,\overline{A})$. Then the claimed result is a direct consequence of the identity between $S_{\ell}$ and $T_{\ell}$ in Lemma˜A.1.
C.15 Proof of Lemma˜6.1
@@
$\overset{\rm(c)}{=}\left[1-\left(1-\frac{\frac{1}{n}\sum_{i=1}^{n}\mathrm{Var}_{P_{i}}(f)}{\mathrm{Var}_{\overline{P}}(f)}\right)^{2}\right]^{-\frac{1}{2}}-1,$
where (a) is due to the data-processing inequality, (b) is the lower-semicontinuity of the $\chi^{2}$ divergence (using arguments similar to [PW25, Theorem 4.9]), and (c) uses
@@
with $u=(u_{1},\dots,u_{n})$ being the unit-length eigenvector of $A$ associated with $\lambda_{2}(A)$. Clearly $\mathbb{E}_{\overline{P}}[f]=\sum_{j=1}^{n}u_{j}=0$ by the orthogonality of $u$ and $\bf 1$ in Lemma˜5.2, and
@@
This shows that $S\geq\lambda_{2}(A)$. [^4] In fact $S=\lambda_{2}(A)$ holds, but we do not need this upper bound.
@@
Here ${\sf LC}(P_{1},P_{2})$ denotes the Le Cam distance, and the last inequality follows from [PW25, Eqn. (7.35)]. Consequently, the second statement follows from
@@
where $I$ and $J$ are the identity and all-ones matrices, respectively. By simple algebra, $A$ is doubly stochastic, and has eigenvalues $\lambda_{1}(A)=1,\lambda_{2}(A)=\cdots=\lambda_{m}(A)=1-\Delta$, and $\lambda_{m+1}(A)=\cdots=\lambda_{mn}(A)=0$. In particular, $\mathrm{Tr}(A)\leq m<1+{\mathsf{C}}$, and the spectral gap of $A$ is $\Delta$.